The First National Bank of Wilson has 680 checking account customers. A recent sample of 40 of these customers showed 27 have a Visa card with the bank.


Construct the 90% confidence interval for the proportion of checking account customers who have a Visa card with the bank. (Use z Distribution Table.) (Round your answers to 3 decimal places.)
Answer #1

= 27 / 40 = 0.675

1 - = 0.325

Critical value at Alpha = 0.10 = 1.645

The Confidence Interval is given by ME, where

The Lower Limit = 0.675 - 0.122 = 0.553

The Upper Limit = 0.675 + 0.122 = 0.797

The 90% Confidence Interval is (0.553 , 0.797)